OpenStudy (anonymous):

Confirm that f and g are inverses by showing that f(g(x)) = x and g(f(x)) = x. f(x) = 9x + 3 and g(x) =x-3/9

OpenStudy (anonymous):

\[g(x) =\frac{x-3 }{ 9}\]

OpenStudy (anonymous):

\[f(g(x))=f(\frac{x-3}{9})=9\times \frac{x-3}{9}+3=x-3+3=x\] for the first one